OpenStudy (anonymous):

Can someone please solve this problem showing me step by step how you get the answer please? -y-1/4x=-3 x being-2

OpenStudy (amistre64):

is that -1/4x or is it (-y-1)/4x?

OpenStudy (amistre64):

\[\frac{-y-1}{4x}=-3?\]

OpenStudy (anonymous):

-1/4x

OpenStudy (amistre64):

\[-y -\frac{1}{4(-2)}=-3\]

OpenStudy (amistre64):

first step is to plug in -2 for x....

OpenStudy (amistre64):

we get: -y + (1/8) = -3 ; -(1/8) from each side to get -y = -3' 1/8 ; /-1 to both sides to get: y = 3' 1/8 = 25/8

OpenStudy (anonymous):

if you -1/8 from both sides how are you getting -3 1/8 ?

OpenStudy (amistre64):

-3 -(1/8) = -3' 1/8

OpenStudy (amistre64):

if you owe someone 3 dollars and you owe the 1/8 more of a dollar the total you owe them is: 3' 1/8 dollars

OpenStudy (anonymous):

OK, got that part..where are you getting the dividing -1 to both sides would it be -y =-1 is that how?

OpenStudy (amistre64):

-y = -1y -1y/-1 = y :)